OpenStudy (anonymous):

Determine pH of 0.165 M NH4Cl

OpenStudy (abb0t):

ammonium chloride is a weak acid, so you'd use: for weak acids use pH = 1/2(pKa - log(c_acid) ) \[pH = \frac{ 1 }{ 2 }(pKa-\log(C_{acid}))\] where C = conjugate acid.